Whats the magnitude of the resultant of the two forces

y 71° F 490 N

F=400 N
Answer #1

Dale FSinti 71 Ecos.74 LE F=4oon 490 S unnation of all Vertical Forces fy= F Cos71 + ugo Ey = 400 Cos71 +490 Ey = 620.22N SumFy R Fox R = 378.2012 +162 2012 +620.22 R=726.435 N

To calculate position of resultant use

Tan (theta) = Fy/Fx

Please rate positively